OpenStudy (cutiecomittee123):

What are the coordinates to the solution that lays in quadrant 1? x^2-y^2=25 x+y=25

OpenStudy (anonymous):

All coordinates in quadrant 1 are positive

OpenStudy (cutiecomittee123):

good to know

OpenStudy (cutiecomittee123):

I tried squaring the second equation and then adding both equations together, which essentially canceled out the y^2's so then I solved for x but came up with an odd number.

OpenStudy (anonymous):

\(x^2-y^2=(x-y)(x+y)=25\) now use the second equation x+y=25 hope this is enough, :)

OpenStudy (cutiecomittee123):

thanks for that but what do I do with that information?

OpenStudy (anonymous):

next step would be: x-y=1 x+y=25

OpenStudy (anonymous):

then 2x=26 x=13 y=12

OpenStudy (cutiecomittee123):

so you basically took the square root of the top equation

OpenStudy (anonymous):

i used the fact that a^2-b^2=(a-b)(a+b)

OpenStudy (cutiecomittee123):

how did 25 go to 1?

OpenStudy (anonymous):

if in (x-y)(x+y)=25 you substitute the value of (x+y) from second equation, you get (x-y)25=25 or x-y=1

OpenStudy (anonymous):

got it?

OpenStudy (cutiecomittee123):

okay

OpenStudy (perl):

Nice solution. Another way to do it is solve for x in the second equation and substitute it in the first equation. That gives you a quadratic and more work involved to get the answer.

OpenStudy (cutiecomittee123):

Yeah I figured it out though:) (13,12)